# NeWRF: A Deep Learning Framework for Wireless Radiation Field Reconstruction and Channel Prediction

### v1

@@ Appendix B Subcarrier Channel Prediction
Modern wireless communication systems, such as WiFi and 5G use Orthogonal Frequency Division Multiplexing (OFDM) for modulation. OFDM divides the spectrum bandwidth into multiple narrow frequency bands, termed subcarriers. Each subcarrier modulates one data symbol. To demodulate data, the receiver has to perform channel estimation for all subcarriers. For instance, a 20 MHz WiFi channel is typically divided into 64 subcarriers, with each subcarrier 312.5 kHz apart from its neighbors. The channels of 52 (out of the 64) subcarriers are typically measured and reported by commercial off-the-shelf WiFi chips Ma et al. [2019], known as channel state information (CSI).
@@
NeWRF is able to predict the subcarrier channels accurately using a model trained only on a single subcarrier without further fine-tuning. To demonstrate this capability, we use a model trained on the conference room dataset with carrier frequency 2.412 GHz to infer the channels for its nearby subcarriers. To synthesize the channel at a specific subcarrier frequency, we simply change the signal frequency $f$ in the channel synthesis algorithm (Eq. (7)) to the subcarrier frequency $f_{sc}=f_{c}+k\times 3.125\times 10^{5}~{}Hz$, where $k$ is the index of the subcarrier, ranging from -26 to 26; and $f_{c}$ is the carrier frequency of the channel. Figure 10(a) shows the predicted amplitude and phase versus the subcarrier indices at five different receiver locations (Pt.1$\sim$ 5). The predictions align very well with the ground truth. Figure 10(b) shows a different view of the same data as 10(a), where we plot the real and imaginary parts of the subcarriers in x and y-axis, respectively.
@@
The reason behind this capability of NeWRF is that the model learns to decouple the attenuation factor and phase rotation due to the reflection from that of free space propagation. The models predicts only the radiated signal, which is nearly unchanged across the subcarrier frequencies. The dominant cause of the varying patterns across subcarriers is the phase shift during the free space propagation and constructive/destructive interference of multipath, which are explicitly modeled by the channel synthesis algorithm.
Appendix C Antenna Array for Direction-of-Arrival Finding
Figure 12: An array of antennas can be used to resolve the signal direction of arrival $\theta$.
In wireless communication, the Direction-of-Arrival (DoA) of a signal can be measured using an array of antennas. Figure 12 shows a 1D antenna array, where each antenna is separated from its neighbor by distance $d$. The signal emitted by a radio transmitter reaches each antenna in the array at slightly different times depending on the signal’s DoA. These time differences lead to phase shifts in the received signal at each antenna, $\Delta\psi$. By examining the phase shift between two antennas, we can resolve the DoA of a signal:
$\theta=\arccos(\frac{\lambda\Delta\psi}{2\pi d}),$
where $\lambda=c/f$ is the signal’s wavelength. Although the phase difference between two antennas is enough to resolve the DoA for a single path, having more antennas in an array can improve the measurement accuracy and resolution. Generally, an antenna array with $N$ antennas can resolve at most $N-1$ paths. In the same manner, a 2D antenna array can be used to measure the DoA in both azimuth and elevation directions.

### v2

@@ Appendix B Subcarrier Channel Prediction
Modern wireless communication systems, such as WiFi and 5G use Orthogonal Frequency Division Multiplexing (OFDM) for modulation. OFDM divides the spectrum bandwidth into multiple narrow frequency bands, termed subcarriers. Each subcarrier modulates one data symbol. To demodulate data, the receiver has to perform channel estimation for all subcarriers. For instance, a 20 MHz WiFi channel is typically divided into 64 subcarriers, with each subcarrier 312.5 kHz apart from its neighbors. The channels of 52 (out of the 64) subcarriers are typically measured and reported by commercial off-the-shelf WiFi chips (Ma et al., 2019), known as channel state information (CSI).
@@
NeWRF is able to predict the subcarrier channels accurately using a model trained only on a single subcarrier without further fine-tuning. To demonstrate this capability, we use a model trained on the conference room dataset with carrier frequency 2.412 GHz to infer the channels for its nearby subcarriers. To synthesize the channel at a specific subcarrier frequency, we simply change the signal frequency $f$ in the channel synthesis algorithm (Eq. (7)) to the subcarrier frequency $f_{sc}=f_{c}+k\times\Delta f~{}Hz$, where $k$ is the index of the subcarrier, ranging from -26 to 26; $\Delta f$ is the subcarrier spacing, which equals to 3.125 kHz for 802.11 a/g/n; and $f_{c}$ is the carrier frequency of the channel. Figure 13(a) shows the predicted amplitude and phase versus the subcarrier indices at five different receiver locations (Pt.0$\sim$ 4). As is shown, the predictions align very well with the ground truth. Figure 13(b) shows a different view of the same data as 13(a), where we plot the real and imaginary parts of the subcarriers in x and y-axis, respectively. In particular, we note that the prediction accuracy remains high even for highly frequency selective channels. For instance, the channel amplitude of Pt.0 varies by orders of magnitude across different subcarriers. And NeWRF is still able to predict the CSI accurately.
@@
The reason behind this capability of NeWRF is that the model learns to decouple the attenuation factor and phase rotation due to the reflection from that of free space propagation. The model predicts only the radiated signal, which is nearly unchanged across the subcarrier frequencies. The dominant cause of the varying patterns across subcarriers is the phase shift during the free space propagation and constructive/destructive interference of multipath, which are explicitly modeled by the channel synthesis algorithm.
Appendix C Impact of Noise
Here, we are interested in seeing the robustness of NeWRF against the influence of noise in wireless measurements. In this experiment, we add artificial Gaussian white noise to the simulated channels. The ratio between the channel power and added noise power is denoted as channel SNR. Figure 13 shows the results, the blue line shows the results of adding a fixed noise, i.e. bias, to the ground truth channel in the training set. As is expected, the prediction accuracy reduces as the channel SNR decreases. However, typical wireless systems today achieve more than 20 dB channel SNR in practice for reliable communication, where NeWRF performs well in such SNRs. Also, we note that collecting multiple measurements with different biases can improve the robustness. The orange line in Figure 13 demonstrates this effect, where we use measurements with different additive noises in each iteration of training. The model is able to maintain good performance even at very low SNRs. This is because the model implicitly averages these measurements to reduce the bias during the training.